OpenStudy (anonymous):

if sin(theta) = 1/3 find the tan(theta) where (theta) terminates in QI

OpenStudy (anonymous):

\[if \sin \Theta =\frac{ 1 }{ 3 } find the \tan \theta where \theta terminates \in QI \]

OpenStudy (solomonzelman):

sin x = 1/3